# LLMs Assist NLP Researchers: Critique Paper (Meta-)Reviewing

### v2

@@ Abstract
This study focuses on the topic of LLMs Assist NLP Researchers, particularly examining the effectiveness of LLM in assisting paper (meta-)reviewing and its recognizability. To address this, we constructed the ReviewCritique dataset, which includes two types of information: (i) NLP papers (initial submissions rather than camera-ready) with both human-written and LLM-generated reviews, and (ii) each review comes with “deficiency” labels and corresponding explanations for individual segments, annotated by experts. Using ReviewCritique, this study explores two threads of research questions: (i) “LLMs as Reviewers”, how do reviews generated by LLMs compare with those written by humans in terms of quality and distinguishability? (ii) “LLMs as Metareviewers”, how effectively can LLMs identify potential issues, such as Deficient or unprofessional review segments, within individual paper reviews? To our knowledge, this is the first work to provide such a comprehensive analysis.
@@
Jiangshu Du1, Yibo Wang1, Wenting Zhao1,10, Zhongfen Deng1, Shuaiqi Liu3, Renze Lou2, Henry Peng Zou1, Pranav Narayanan Venkit2, Nan Zhang2, Mukund Srinath2, Haoran Ranran Zhang2, Vipul Gupta2, Yinghui Li4, Tao Li5, Fei Wang6, Qin Liu7, Tianlin Liu8, Pengzhi Gao9, Congying Xia10, Chen Xing11, Jiayang Cheng12, Zhaowei Wang12, Ying Su12, Raj Sanjay Shah13, Ruohao Guo13, Jing Gu14, Haoran Li15, Kangda Wei16, Zihao Wang12, Lu Cheng1, Surangika Ranathunga17, Meng Fang18, Jie Fu12, Fei Liu21, Ruihong Huang16, Eduardo Blanco19, Yixin Cao20, Rui Zhang2, Philip S. Yu1, Wenpeng Yin2 1University of Illinois Chicago, 2Penn State University, 3Hong Kong Polytechnic University, 4Tsinghua Univeristy, 5University of Utah, 6University of Southern California, 7University of California, Davis, 8University of Basel, Switzerland, 9Xiaomi AI Lab, 10Salesforce Research, 11Scale AI, 12HKUST, 13Georgia Institute of Technology, 14University of California, Santa Cruz, 15Singapore University of Technology and Design, 16Texas A&M University, 17Massey University, New Zealand, 18University of Liverpool, 19University of Arizona, 20Fudan University, 21Emory University jdu25@uic.edu, wenpeng@psu.edu
@@ 2 Related Work
Researchers have explored various aspects of AI for reviews. One area of interest is the use of AI to assist in automatically generating peer reviews, such as predicting scores Li et al. (2020); Zhou et al. (2024); Wang et al. (2020) and writing reviews Gao et al. (2024); Wang et al. (2020); Yuan et al. (2022); Liu and Shah (2023) and meta-reviews Li et al. (2023); Lin et al. (2023). Another line of research focuses on leveraging NLP methods to evaluate the quality of human reviews Xiong and Litman (2011); Guo et al. (2023); Kumar et al. (2023); Ghosal et al. (2022b).
To facilitate research on AI for peer review, several datasets have been introduced. PeerRead Kang et al. (2018), MOPRD Lin et al. (2023), and NLPeer Dycke et al. (2023) are datasets containing a large number of peer reviews and their corresponding papers but without expert annotations. Other datasets focus on specific aspects of peer reviews, such as argument Kennard et al. (2022); Hua et al. (2019); Yuan et al. (2022); Cheng et al. (2020); Ruggeri et al. (2023), politeness Bharti et al. (2023), uncertainty detection Ghosal et al. (2022b), contradictions in review pairs Kumar et al. (2023), and substantiation Guo et al. (2023). Peer Review Analyze Ghosal et al. (2022a) annotates reviews across four facets: paper section correspondence, aspect, functionality, and significance. However, these datasets are solely based on reviews and none of them are highly expert-demanding. In contrast, ReviewCritique  is the first dataset to benchmark LLMs’ capability as a responsible meta-reviewer.
Recently, researchers have also explored the evaluation of LLMs’ deficiency and limitations in automatic paper reviewing tasks Zhou et al. (2024); Liu and Shah (2023); Robertson (2023); Liang et al. (2023). Our work differs from previous works in that we provide a quantitative comparison of human-written and LLM-generated paper reviews at the sentence level. This fine-grained analysis allows us to identify specific areas where LLMs excel or struggle in generating high-quality reviews. We also propose a novel metric to measure LLM-generated review diversity.
@@ 3.1 Paper Submission & Review Collection
#Segments 11,376 6,027 5,349 1,612 798 814
    Deficient 713 317 396 221 106 115
    Deficient  pct. (%) 6.27 5.26 7.40 13.71 13.28 14.13
#ExplainationTokens 14773 6957 7816 4156 1978 2178
@@ 3.3 Data Statistics
Table 1 provides the statistics for our ReviewCritique dataset. It shows that 71.57% of human-written reviews and 100% of LLM-generated reviews have Deficient segments. A comparison of accepted and rejected submissions reveals that rejected papers consistently contain a higher percentage of Deficient segments in both human-written and LLM-generated reviews.
@@ 4.1.1 Error type analysis for deficiency
Misunderstanding 22.86 10.41
Neglect 19.64 4.52
Inexpert Statement 18.23 5.88
@@
Out-of-scope 4.35 31.67
Misunderstanding 22.86 10.41
Superficial Review 2.66 9.95
@@
Compared to humans, LLMs are more likely to suggest out-of-scope experiments or analyses. They make significantly fewer "Inexpert Statement" errors. Based on our observations, this because their reviews are usually paper-unspecific and superficial, avoiding expert-level mistakes. Additionally, LLM-generated reviews do not exhibit errors like "Missing Reference," "Invalid Reference," and "Concurrent Work" since they do not point to specific works or provide references.
@@ 4.1.2 Fine-grained review analysis
The Summary section in LLM-generated reviews exhibits relatively better quality compared to other aspects. Our annotators identified only 1.36% of segments as "Inaccurate Summary" among all LLM Deficient segments, which constitutes 0.19% of all LLM-generated segments. In comparison, 5.75% of segments were identified as "Inaccurate Summary" among all Deficient segments in human-written reviews, accounting for 0.36% of all human-written review segments. This is nearly twice the percentage found in LLM-generated summaries. Moreover, error types such as “Summary Too Short” and “Copy-pasted Summary”, which are present in human reviews, were not observed in LLM-generated reviews, suggesting that LLMs are capable of generating summaries of satisfying quality and avoid directly copying content from the paper.
@@
The most dominant type of Deficient in LLM reviews is “Out-of-scope”, accounting for 31.67% of all Deficient segments in LLM-generated reviews (see Table 3). LLMs often highlight weaknesses such as the need for more experiments, lack of generalizability, additional tasks, more analysis, evaluation on languages beyond English, etc. While occasionally relevant, these suggestions often fall outside the paper’s scope and shouldn’t be considered weaknesses.
@@
In addition to generating reviews, we asked LLMs to rate each paper on a scale of 1-10, matching the ICLR and NeurIPS system, for directly comparison with human reviewers. Experiment shows that LLMs tend to give high scores to all submissions, regardless of quality or acceptance status, with averages of 7.43 for accepted and 7.47 for rejected papers. In contrast, human reviewers differentiate more effectively, with averages of 6.41 for accepted and 4.81 for rejected submissions. Thus, LLMs fail to distinguish between accepted and rejected papers, assigning similarly high scores to both.
@@ 4.1.3 Review Diversity
$\vspace{-2mm}O_{i}^{j}=\sum_{k=1}^{n_{j}}\mathbb{I}(\text{sim}(s_{i}^{j},s_{k}^{j})\geq t)\cdot\text{sim}(s_{i}^{j},s_{k}^{j}),\vspace{-1mm}$ (2)
@@
$\vspace{-2mm}R_{i}^{j}=\sum_{l=1}^{m}\mathbb{I}\left(\max_{p}\text{sim}(s_{i}^{j},s_{p}^{l})\geq t\right)\cdot\max_{p}\text{sim}(s_{i}^{j},s_{p}^{l}),$ (3)
@@ 4.2 LLMs as Metareviewers
To mitigate the impact of prompt-specific performance, we employ two prompting strategies: 1) Labeling-All: Given everything necessary including a list of indexed review segments, require the LLM to output a list of triples like (id, Deficient or not, explanation); 2) Select-Deficient: Given everything necessary including a list of indexed review segments, require the LLM to output a list of tuples, (id, explanation), when it believes the “id” corresponds to an Deficient segment.
The detailed prompt templates are in Table 12 and 13 (Appendix F).
@@
The results in Table 5 show that overall scores for all LLMs are relatively low, indicating they can identify some Deficient segments but struggle to articulate their reasoning.
Among the LLMs, Claude Opus achieves the highest scores across all metrics, suggesting its explanations align best with human annotators.
Claude Opus also excels in identifying Deficient segments, as shown previously. GPT-4 and Gemini 1.5 show similar performance to Claude Opus. The open-source models, Llama3 (8B and 70B) and Qwen2-72B, generally score lower than the closed-source models.
@@ 5 Conclusion
While our work provides valuable insights into the potential of LLMs in the peer review process, there are some limitations to consider. During the evaluation of LLMs, ReviewCritique primarily focuses on the textual information from the submissions and does not include figures, tables, or other visual elements. Incorporating these additional components could provide a more comprehensive assessment of LLMs’ capabilities in the peer review process. Additionally, the dataset is currently limited to the NLP domain. It would be interesting to explore the performance of LLMs in other research areas. Expanding the dataset to include papers from various domains could help assess the generalizability of our findings and identify potential domain-specific challenges. Furthermore, our work focuses on the pre-rebuttal phase of the peer review process, assessing reviews based on the original submission. Incorporating the multi-turn aspect of peer review, including author rebuttals and post-rebuttal reviews, could offer a more comprehensive understanding of LLMs’ capabilities in the entire review process.
@@ Appendix D Deficient Segment Error Types
Out-of-scope 4.35% 5.05% 3.79% 31.67% 34.91% 28.70%
Inaccurate Summary 5.75% 8.52% 3.54% 1.36% 0.94% 1.74%
Neglect 19.64% 24.29% 15.91% 4.52% 5.66% 3.48%
Inexpert Statement 18.23% 16.72% 19.44% 5.88% 4.72% 6.96%
Misunderstanding 22.86% 17.35% 27.27% 10.41% 9.43% 11.30%
Vague Critique 9.26% 5.99% 11.87% 7.69% 7.55% 7.83%
Misinterpret Novelty 3.79% 6.94% 1.26% 1.36% 1.89% 0.87%
@@
Writing 2.81% 2.52% 3.03% 4.07% 2.83% 5.22%
Superficial Review 2.66% 3.15% 2.27% 9.95% 11.32% 8.70%

### v3

@@ Abstract
This study focuses on the topic of LLMs Assist NLP Researchers, particularly examining the effectiveness of LLM in assisting paper (meta-)reviewing and its recognizability. To address this, we constructed the ReviewCritique dataset, which includes two types of information: (i) NLP papers (initial submissions rather than camera-ready) with both human-written and LLM-generated reviews, and (ii) each review comes with “deficiency” labels and corresponding explanations for individual segments, annotated by experts. Using ReviewCritique, this study explores two threads of research questions: (i) “LLMs as Reviewers”, how do reviews generated by LLMs compare with those written by humans in terms of quality and distinguishability? (ii) “LLMs as Metareviewers”, how effectively can LLMs identify potential issues, such as Deficient or unprofessional review segments, within individual paper reviews? To our knowledge, this is the first work to provide such a comprehensive analysis. Our dataset is available at https://github.com/jiangshdd/ReviewCritique.
@@
Jiangshu Du1, Yibo Wang1, Wenting Zhao1,10, Zhongfen Deng1, Shuaiqi Liu3, Renze Lou2, Henry Peng Zou1, Pranav Narayanan Venkit2, Nan Zhang2, Mukund Srinath2, Ranran Haoran Zhang2, Vipul Gupta2, Yinghui Li4, Tao Li5, Fei Wang6, Qin Liu7, Tianlin Liu8, Pengzhi Gao9, Congying Xia10, Chen Xing11, Cheng Jiayang12, Zhaowei Wang12, Ying Su12, Raj Sanjay Shah13, Ruohao Guo13, Jing Gu14, Haoran Li15, Kangda Wei16, Zihao Wang12, Lu Cheng1, Surangika Ranathunga17, Meng Fang18, Jie Fu12, Fei Liu21, Ruihong Huang16, Eduardo Blanco19, Yixin Cao20, Rui Zhang2, Philip S. Yu1, Wenpeng Yin2 1University of Illinois Chicago, 2Penn State University, 3Hong Kong Polytechnic University, 4Tsinghua Univeristy, 5Google DeepMind, 6University of Southern California, 7University of California, Davis, 8University of Basel, Switzerland, 9Xiaomi AI Lab, 10Salesforce Research, 11Scale AI, 12HKUST, 13Georgia Institute of Technology, 14University of California, Santa Cruz, 15Singapore University of Technology and Design, 16Texas A&M University, 17Massey University, New Zealand, 18University of Liverpool, 19University of Arizona, 20Fudan University, 21Emory University jdu25@uic.edu, wenpeng@psu.edu
@@ 2 Related Work
Researchers have explored various aspects of AI for reviews. One area of interest is the use of AI to assist in automatically generating peer reviews, such as predicting scores Li et al. (2020); Zhou et al. (2024); Wang et al. (2020); Deng et al. (2020) and writing reviews Gao et al. (2024); Wang et al. (2020); Yuan et al. (2022); Liu and Shah (2023) and meta-reviews Li et al. (2023); Lin et al. (2023b). Another line of research focuses on leveraging NLP methods to evaluate the quality of human reviews Xiong and Litman (2011); Guo et al. (2023); Kumar et al. (2023); Ghosal et al. (2022b).
To facilitate research on AI for peer review, several datasets have been introduced. PeerRead Kang et al. (2018), MOPRD Lin et al. (2023b), and NLPeer Dycke et al. (2023) are datasets containing a large number of peer reviews and their corresponding papers but without expert annotations. Other datasets focus on specific aspects of peer reviews, such as argument Kennard et al. (2022); Hua et al. (2019); Yuan et al. (2022); Cheng et al. (2020); Ruggeri et al. (2023), politeness Bharti et al. (2023), uncertainty detection Ghosal et al. (2022b), contradictions in review pairs Kumar et al. (2023), and substantiation Guo et al. (2023). Peer Review Analyze Ghosal et al. (2022a) annotates reviews across four facets: paper section correspondence, aspect, functionality, and significance. However, these datasets are solely based on reviews and none of them are highly expert-demanding. In contrast, ReviewCritique  is the first dataset to benchmark LLMs’ capability as a responsible meta-reviewer.
Recently, researchers have also explored the evaluation of LLMs’ deficiency and limitations in automatic paper reviewing tasks Zhou et al. (2024); Liu and Shah (2023); Robertson (2023); Liang et al. (2023); Lin et al. (2023a). Our work differs from previous works in that we provide a quantitative comparison of human-written and LLM-generated paper reviews at the sentence level. This fine-grained analysis allows us to identify specific areas where LLMs excel or struggle in generating high-quality reviews. We also propose a novel metric to measure LLM-generated review diversity.
@@ 3.1 Paper Submission & Review Collection
#Segments 11,376 6,027 5,349 1,611 812 799
    Deficient 713 317 396 225 144 81
    Deficient  pct. (%) 6.27 5.26 7.40 13.97 17.73 10.14
#ExplainationTokens 14,773 6,957 7,816 3,877 2,584 1,293
@@ 3.3 Data Statistics
Table 1 provides the statistics for our ReviewCritique dataset. We compare from two dimensions. First, at both the review and segment granularity, LLM-generated reviews contain more Deficient instances compared to human-written reviews (100% vs. 71.57% at the review level, and 13.97% vs. 6.27% at the segment level). Next, we compared “Accepted” and “Rejected”, which generally correspond to “higher-quality” and “lower-quality” submissions, respectively. Notably, LLM-generated reviews demonstrated a higher frequency of Deficient segments for “Accepted” submissions than for “Rejected” ones, which contrasts with what we observed in human-written reviews. Drawing from our analysis in the “Weaknesses” part of Section 4.1.2, we suggest the following explanation: human reviewers are often able to sense the overall quality of a paper. If they believe a submission is of poor quality and intend to reject it, they tend to collect more weaknesses to justify their decision, which sometimes leads to overemphasis on the paper’s flaws. In contrast, LLMs lack the ability to discern paper quality and often generate superficial and non-specific criticisms equally for both “Accepted” and “Rejected” submissions, making these review segments more likely to be inaccurate for higher-quality papers.
@@ 4.1.1 Error type analysis for deficiency
Misunderstanding 22.86 9.87
Neglect 19.64 5.83
Inexpert Statement 18.23 6.73
@@
Out-of-scope 4.35 30.49
Misunderstanding 22.86 9.87
Superficial Review 2.66 9.42
@@
Compared to humans, LLMs are more likely to suggest out-of-scope experiments or analyses. They make significantly fewer "Inexpert Statement" errors. Based on our observations, this is because their reviews are usually paper-unspecific and superficial, avoiding expert-level mistakes. Additionally, LLM-generated reviews do not exhibit errors like "Missing Reference," "Invalid Reference," and "Concurrent Work" since they do not point to specific works or provide references.
@@ 4.1.2 Fine-grained review analysis
The Summary section in LLM-generated reviews exhibits relatively better quality compared to other aspects. Our annotators identified only 1.35% of segments as "Inaccurate Summary" among all LLM Deficient segments, which constitutes 0.19% of all LLM-generated segments. In comparison, 5.75% of segments were identified as "Inaccurate Summary" among all Deficient segments in human-written reviews, accounting for 0.36% of all human-written review segments. This is nearly twice the percentage found in LLM-generated summaries. Moreover, error types such as “Summary Too Short” and “Copy-pasted Summary”, which are present in human reviews, were not observed in LLM-generated reviews, suggesting that LLMs are capable of generating summaries of satisfying quality and avoid directly copying content from the paper.
@@
The most dominant type of Deficient in LLM reviews is “Out-of-scope”, accounting for 30.49% of all Deficient segments in LLM-generated reviews (see Table 3). LLMs often highlight weaknesses such as the need for more experiments, lack of generalizability, additional tasks, more analysis, evaluation on languages beyond English, etc. While occasionally relevant, these suggestions often fall outside the paper’s scope and shouldn’t be considered weaknesses.
@@
In addition to generating reviews, we asked LLMs to rate each paper on a scale of 1-10, matching the ICLR and NeurIPS system, for directly comparison with human reviewers. Experiment shows that LLMs tend to give high scores to all submissions, regardless of quality or acceptance status, with averages of 7.43 for accepted and 7.47 for rejected papers. In contrast, human reviewers scored averages of 6.41 for accepted and 4.81 for rejected submissions. Thus, LLMs do not align with the human reviewers in discerning paper quality based on their internal scoring mechanism.
@@ 4.1.3 Review Diversity
$O_{i}^{j}=\sum_{k=1}^{n_{j}}\mathbb{I}(\text{sim}(s_{i}^{j},s_{k}^{j})\geq t)\cdot\text{sim}(s_{i}^{j},s_{k}^{j}),$ (2)
@@
$R_{i}^{j}=\sum_{l=1}^{m}\mathbb{I}\left(\max_{p}\text{sim}(s_{i}^{j},s_{p}^{l})\geq t\right)\cdot\max_{p}\text{sim}(s_{i}^{j},s_{p}^{l}),$ (3)
@@ 4.2 LLMs as Metareviewers
To mitigate the impact of prompt-specific performance, we employ two prompting strategies: 1) Labeling-All: Given everything necessary including a list of indexed review segments, require the LLM to output a list of triples like (id, Deficient or not, explanation); 2) Select-Deficient: Given everything necessary including a list of indexed review segments, require the LLM to output a list of tuples, (id, explanation), when it believes the “id” corresponds to an Deficient segment. The detailed prompt templates are in Table 12 and 13 (Appendix F).
@@
The results in Table 5 show that overall scores for all LLMs are relatively low, indicating they can identify some Deficient segments but struggle to articulate their reasoning. Among the LLMs, Claude Opus achieves the highest scores across all metrics, suggesting its explanations align best with human annotators. Claude Opus also excels in identifying Deficient segments, as shown previously. GPT-4 and Gemini 1.5 show similar performance to Claude Opus. The open-source models, Llama3 (8B and 70B) and Qwen2-72B, generally score lower than the closed-source models.
@@ 5 Conclusion
While our work provides valuable insights into the potential of LLMs in the peer review process, there are some limitations. During the evaluation of LLMs, ReviewCritique primarily focuses on the textual information from the submissions and does not include figures, tables, or other visual elements. Incorporating these additional components could provide a more comprehensive assessment of LLMs’ capabilities in the peer review process. Additionally, the dataset is currently limited to the NLP domain. It would be interesting to explore the performance of LLMs in other research areas. Expanding the dataset to include papers from various domains could help assess the generalizability of our findings and identify potential domain-specific challenges. Furthermore, our work focuses on the pre-rebuttal phase of the peer review process, assessing reviews based on the original submission. Incorporating the multi-turn aspect of peer review, including author rebuttals and post-rebuttal reviews, could offer a more comprehensive understanding of LLMs’ capabilities in the entire review process.
Ethical Considerations
In this study, we carefully considered the ethical implications of using LLMs to assist in the review process. We acknowledge potential risks, including bias, lack of accountability, and the possibility of undermining the integrity of scientific evaluations. Importantly, this work does not advocate for the use of LLMs in paper (meta-)reviewing. Rather, our study underscores that LLMs are currently insufficient to replace human reviewers, especially in tasks that require expert judgment and nuanced understanding.
Acknowledgments
The authors appreciate the reviewers for their insightful comments and suggestions. This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
@@ Appendix D Deficient Segment Error Types
Out-of-scope 4.35% 5.05% 3.79% 30.49% 33.80% 24.69%
Inaccurate Summary 5.75% 8.52% 3.54% 1.35% 0.70% 2.47%
Neglect 19.64% 24.29% 15.91% 5.83% 4.23% 8.64%
Inexpert Statement 18.23% 16.72% 19.44% 6.73% 4.23% 11.11%
Misunderstanding 22.86% 17.35% 27.27% 9.87% 11.97% 6.17%
Vague Critique 9.26% 5.99% 11.87% 7.17% 4.23% 12.35%
Misinterpret Novelty 3.79% 6.94% 1.26% 2.24% 2.82% 1.23%
@@
Writing 2.81% 2.52% 3.03% 4.48% 4.23% 4.94%
Superficial Review 2.66% 3.15% 2.27% 9.42% 7.04% 13.58%